OpenStudy (anonymous):

A bug walking on the circular rim of a flower pot completes one lap around the planter in 12.3 s. If the radius of the planter is 11.5 cm, how fast was it traveling? Answer in units of cm/s How fast in m/s? Answer in units of m/s

OpenStudy (anonymous):

5.871

OpenStudy (anonymous):

circumference of circle = 2*pi*11.5 cm speed of the bug = (2*pi*11.5) / ( 12.3) cms / sec = 5.87 cm/s = .0587 m/s

OpenStudy (anonymous):

That's a big planter.